Find an equation for the line parallel to
y=10x+4
and goes through the point
(6,8)
.
Write your answer in the form
y
=
m
x
+
b
y=mx+b
.

Found 3 solutions by josgarithmetic, addingup, MathTherapy:
Answer by josgarithmetic(39617)
Answer by addingup(3677)

y = mx+b
y = 10x+4
set x to 0:
y = 10(0) + 4
.
10*0 = 0, so:
y = 4
OK, we've got one of our numbers. When x = 0 y = 4. We write this: (0, 4) and that's our first pair.
.
Now let's rewrite the equation and set y = 0:
10x = -y + 4
x = (-y+4)/10
x = -0+4/10
-(0 + 4/10) makes -0.4.
Now we can write the other pair: (-0.4, 0)
These are called ordered pairs and can now be plotted. See plot below, draw a straight line joining the two points.
.
Parallel line Two lines are parallel if they have the same slope.
We know our slope is 10 and our y-intercept you say is 8. Below is the graph showing parallel lines:





Answer by MathTherapy(10552)

Find an equation for the line parallel to
y=10x+4
and goes through the point
(6,8)
.
Write your answer in the form
y
=
m
x
+
b
y=mx+b
.
I have a question that I hope will be answered. 
What the H..L does @addingup's answer have to do with the question asked?
Wasn't it JUST an EQUATION thatw's requested?
In addition, was a GRAPH requested? I still don't know what some of these people are doing in this forum. I guess they're FAKE WANT-TO-BE math aides.